OpenStudy (anonymous):

y = 55 x. Use the x values for the number of hours: 0, 10, and 20. A:{(0, 0), (550, 10), (1100, 20)} B:{(0, 0), (10, 550), (20, 1010)} C:{(0, 0), (10, 550), (20, 1100)} D: {(0, 0), (10, 55), (1100, 20)}

OpenStudy (anonymous):

C:{(0, 0), (10, 550), (20, 1100)}

OpenStudy (phi):

this question is asking for the (x,y) pairs you get from y= 55x (y is 55 times x) when x= 0, 10 and 20 you get three pairs replace x with 0 in the equation y=55 x to find y= 55*0= 0 so (0,0) is one of the pairs now replace x with 10: y= 55*10= 550, so (10,550) is a pair finally, replace x with 20: y= 55*20= 1100 so (20,1100) is the last pair